OpenStudy (anonymous):

what's the value of g(0)?

OpenStudy (anonymous):

what is g(x)?

OpenStudy (anonymous):

the problem says .... f(x)=2x+5 and g(x)= -1 over 3x +2 .. find each value

OpenStudy (anonymous):

so 20. it says g(0) fid the value .

OpenStudy (anonymous):

plug 0 in place of x then simplify

OpenStudy (anonymous):

ugh , i hate math .. thats why im on here so people can ive me the answer and tell me how they got it !

OpenStudy (anonymous):

g(0)= -1/(3(0)+2))

OpenStudy (anonymous):

g(0)= -1/2

OpenStudy (anonymous):

Thank you so much . . what is the value for f(1) ?

OpenStudy (anonymous):

plug 1 in place of x in f(x) f(1)=2(1)+5 f(1)=2+5 f(1)=7

OpenStudy (anonymous):

what is he value of f(-4) ?

OpenStudy (anonymous):

for this one u try plugging in (-4) to f(x).

OpenStudy (anonymous):

ugh , ok i guess ima try it .

OpenStudy (anonymous):

i fgot 10 , is that corrct?

OpenStudy (anonymous):

Is my answer correct ?

OpenStudy (anonymous):

-8 + 5= -3